OpenStudy (anonymous):

Question #1 Michele is making two different recipes for nut bread. One calls for 1/8 cup of nuts. The other calls for 3/4 cup of nuts. How many cups of nuts will Michele need in all? Question #2 How many cups of nuts would Michele need is she doubled the recipe?

OpenStudy (anonymous):

Q1 3/4=6/8 6/8+1/8=7/8 cups of nuts needed total. Q2 2*7/8= 14/8 or 7/4

OpenStudy (anonymous):

Thank you so much! 😀

OpenStudy (anonymous):

cheers

OpenStudy (anonymous):

ans1: 1/8+3/4=13/8 ans2: 2*13/8=13/4